OpenStudy (anonymous):

Solve the equation -2(5 - 3n) + 8 = 4 + 5n for n

OpenStudy (anonymous):

OK, so first you want to distribute out the negative 2. What do you get?

OpenStudy (anonymous):

-2(5-3n)+8=4+5n -10+6n+8=4+5n -2+6n=4+5n 6n-5n=4+2 n=6